• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 11:07:39 on March 6,2007

    Entry number 194943

    keyword=left ioc reboot, control computer frozen

    left iof got stuck, rebooted, but it not shown it's coming back
    on the hacsbc2 screen. The medm screen frozen. Had to logout and re-login in. The control screen is fine now. Noticed left NMR lost lock during the process and dipole field drifted away. Stopped the
    run 1205, which had dipole field drift near the end of run.
    All is fine now. Will need to re-start all the monitoring charts.

    FIGURE 1